# On the eigenvalue problem for a bulk/surface elliptic system

### v1

@@ Abstract.
where $\Omega$ is a bounded open subset of $\mathbb{R}^{N}$ ($N\geq 2$) with $C^{1}$ boundary $\Gamma=\Gamma_{0}\cup\Gamma_{1}$, $\Gamma_{0}\cap\Gamma_{1}=\emptyset$, $\Gamma_{1}$ being nonempty and relatively open on $\Gamma$. Moreover $\mathcal{H}^{N-1}(\overline{\Gamma}_{0}\cap\overline{\Gamma}_{1})=0$ and $\mathcal{H}^{N-1}(\Gamma_{0})>0$. We recognize that $L^{2}(\Omega)\times L^{2}(\Gamma_{1})$ admits a Hilbert basis of eigenfunctions of the problem and we describe the eigenvalues. Moreover, when $\Gamma$ is at least $C^{2}$ and $\overline{\Gamma}_{0}\cap\overline{\Gamma}_{1}=\emptyset$, we give several qualitative properties of the eigenfunctions.
@@ 1.1. Presentation of the problem and literature overwiew
where $\Omega$ is a bounded open subset of $\mathbb{R}^{N}$ ($N\geq 2$) with $C^{r}$ boundary $\Gamma$ (see [21]), with $r=1,2,\ldots,\infty$, $r=1$ when nothing is said. We also assume that $\Gamma=\Gamma_{0}\cup\Gamma_{1}$, $\Gamma_{0}\cap\Gamma_{1}=\emptyset$, $\Gamma_{1}$ being nonempty and relatively open on $\Gamma$ (or equivalently $\overline{\Gamma_{0}}=\Gamma_{0}$). Denoting by $\mathcal{H}^{N-1}$ the Hausdorff measure, we also assume that $\mathcal{H}^{N-1}(\overline{\Gamma}_{0}\cap\overline{\Gamma}_{1})=0$ and $\mathcal{H}^{N-1}(\Gamma_{0})>0$. These properties of $\Omega$, $\Gamma_{0}$ and $\Gamma_{1}$ will be assumed, without further comments, throughout the paper.
In problem (1) $\lambda$ is a real or complex parameter and we respectively denote by $\Delta$ and $\Delta_{\Gamma}$ the Laplace operator in $\Omega$ and the Laplace–Beltrami operator on $\Gamma$, while $\nu$ stands for the outward normal to $\Omega$. We shall look for eigenvalues and eigenfunctions on problem (1), that is for values of $\lambda$ for which (1) has a nontrivial (real or complex–valued) solution, i.e. an eigenfunction.
Problem (1) has been studied, as a particular case, in [27], when $\Gamma_{0}=\emptyset$ and $\lambda,u$ are real. The authors refer to this case as to the Dirichlet one. The study in [27] is motivated by several papers on the Allen–Cahn equation subject to a dynamic boundary condition, see [7, 9, 40]. Indeed finding a Hilbert basis of eigenfunctions allows to look for solutions of the evolution problem by using a Faedo–Galerkin scheme.
@@
As to the author’s knowledge the case $\Gamma_{0}\not=\emptyset$ in problem (1) has not yet been considered in the mathematical literature. The motivation for studying it originates from a different evolution problem, that is the wave equation with hyperbolic boundary conditions. To describe it we introduce the evolutionary boundary value problem
@@
where $\Omega$, $\Gamma_{0}$ and $\Gamma_{1}$ are as above, $w=w(t,x)$, $t\in\mathbb{R}$, $x\in\Omega$, $\Delta=\Delta_{x}$ and $\Delta_{\Gamma}$ denote the Laplacian and Laplace–Beltrami operators with respect to the space variable.
@@
where $u$ is nontrivial and real valued. The function $w$ defined in (3) solves, at least formally, problem (2) if and only if $u$ solves problem (1) with $\lambda=\omega^{2}>0$.
Hence, in the analysis of problem (2), a deep understanding of the eigenvalue problem (1) would allow to find solutions by the Fourier method, provided one can find a complete system of eigenfunctions. Since in the analysis of problem (2) it is usefull to consider complex valued solutions, in the sequel we shall consider complex–valued functions everywhere. By the way the due attention will be given to the real case.
Since our study of problem (1) is motivated by problem (2), it is worth to give a brief overview of the literature on it. Problems with hyperbolic or, more generally, kinetic boundary conditions, arise in several physical applications. A one dimensional model was studied by several authors to describe transversal small oscillations of an elastic rod with a tip mass on one endpoint, while the other one is pinched. See [2, 10, 12, 22, 33].
A two dimensional model, introduced in [20], more closely motivates problem (2). Hence we shall briefly describe it. One considers a vibrating membrane of surface density $\mu>0$, subject to a tension $T>0$, both taken constant and normalized for simplicity. Moreover, $w=w(t,x)$, $t\in\mathbb{R}$, $x\in\Omega\subset\mathbb{R}^{2}$, denotes the vertical displacement from the rest state. After a standard linear approximation $w$ satisfies the wave equation $w_{tt}-\Delta w=0$ in ×$\mathbb{R}\times\Omega$. Now one supposes that a part $\Gamma_{0}$ of the boundary is pinched, while the other part $\Gamma_{1}$ carries a linear mass density $m>0$ and it is subject to a linear tension $\tau>0$, both taken constant for simplicity.
A practical example of this situation is given by a drumhead with a hole in the interior having a thick border. This situation occurs in bass drums, as one can realize by looking at several pictures of bass drums in the internet. During the paper we shall constantly refer to this motivating example as to the bass drum.
@@
We also would like to point out that, wehn $\Gamma_{0}=\emptyset$ and $\Omega=\mathbb{R}^{N}_{+}$, problem (2) also shows up in Quantum Field Theory, see [53].
Several papers in the literature address the wave equation with kinetic boundary conditions. This fact is even more evident if one takes into account that, plugging the wave equation in (2) into the boundary condition, we can rewrite it as $-\Delta_{\Gamma}w+\partial_{\nu}w+\Delta w=0$, that is a generalized Wentzell (also spelled Ventcel) boundary condition. We refer to [15, 18, 34, 35, 51, 52], and also to the series of papers [46, 47, 48, 49] by the author. All the quoted papers deal with well–posedness issues for variously, linearly or nonlinearly, perturbed versions of (2).
@@
The analysis of the literature made above shows that the simple unperturbed problem (2), which well–posedness is nowadays rather standard, see for example [47], was object of a detailed study only in the case $\Gamma_{0}=\emptyset$ and $\Omega=\mathbb{R}^{N}_{+}$, see [53].
The aim of the present paper is to start such a study. In particular we shall analyze the eigenvalue problem (1). Our aim is to show that most of the classical results concerning the homogeneous Dirichlet problem for the Helmholtz equation, that is the eigenvalue problem
@@
continue to hold for problem (1), in a suitably modified form. The parallelism will go much further than one can expect. To illustrate the last assertion we shall present our main results in the sequel. Preliminarily we are now going to introduce some basic notation.
@@ 1.2. Functional spaces
where $\nabla_{\Gamma}$ denotes the Riemannian gradient on $\Gamma$, and $(\cdot,\cdot)|_{\Gamma}$ the unique Hermitian extension to the tangent bundle of the Riemannian metric on $\Gamma$, see § 2.2. We also denote $|\cdot|_{\Gamma}=(\cdot,\cdot)|_{\Gamma}^{1/2}$.
@@ 1.3. Main results
Moreover, an eigenvalue for (1) will be $\lambda\in\mathbb{C}$ for which (1) has a nontrivial weak solution, which is called an eigenfunction. Finally, for each eigenvalue $\lambda$, the subspace of weak solutions of (1) will be called eigenspace, and its dimension the eigenvalue’s multiplicity.
@@
Problem (1) has countably many eigenvalues, all of which are real and of finite multiplicity, constituting the set $\Lambda\subset\mathbb{R}$. By repeating each eigenvalue according to its multiplicity we can write $\Lambda=\{\lambda_{n},n\in\mathbb{N}\}$, where
(11) $0<\lambda_{1}\leq\lambda_{2}\leq\cdot\leq\lambda_{n}\leq\lambda_{n+1}\leq\cdots,\qquad\lambda_{n}\to\infty\quad\text{as $n\to\infty$.}$
Moreover there exists an orthogonal sequence $(u_{n})_{n}$ in $H^{1}$ such that each $u_{n}$ is a real–valued eigenfunction corresponding to $\lambda_{n}$, i.e. is a nontrivial weak solution of
@@
Theorem 1 generalizes [27, Theorem 4.4], in the Dirichlet case $K=0$ and $\alpha=1$ to $\Gamma_{0}\not=\emptyset$ and to the complex case.
@@
It extends to problem (1) well–known results concerning problem (4), see for example [17] and [14]. The proof is inspired by the quoted references.
@@
Indeed we shall consider the case in which $\Gamma_{0}$ and $\Gamma_{1}$ are both relatively open on $\Gamma$, so $\Gamma$ has at least two connected components. We remark that, in the example of the bass drum described above, $\Gamma_{0}$ and $\Gamma_{1}$ are exactly these two connected components, so this assumption is rather natural. Moreover, we shall also assume $\Gamma$ to be at least $C^{2}$.
To avoid repeating these assumptions several times we formalize them as follows
@@
both endowed with the norms inherited from the products. Our third main result concerns the regularity of the eigenfunctions in Theorem 1. In addition to its independent interest, it is an essential tool in pursuing the goal described above.
@@
Let assumption (R) holds and for all $n\in\mathbb{N}$ let $u_{n}$ be a weak solution of (12). Then each $u_{n}$ satisfies the following further regularity properties:
@@
The proof of Theorem 3 is based on appropriate regularity results for problem (13), see § 3 below, and on a bootstrap argument.
Our last main result brings the parallelism between problems (1) and (4) to an unexpected level. Indeed, despite of the difference between the two problems, their principal eigenspaces and eigenfunctions exhibit similar properties.
@@
Let assumption $R$ holds and $\Omega$ be connected. Then the principal eigenvalue $\lambda_{1}$ in Theorem 1 is simple, i.e. $\lambda_{1}<\lambda_{2}$ in (11), and $u_{1}$ has constant sign in $\Omega\cup\Gamma_{1}$.
Remark 2.
The most unexpect part of the statement is probably that $u_{1}$ has constant sign also on $\Gamma_{1}$.
Finally, we would like to point out an interesting consequence of our main results in the radial case. Let $\Omega=B_{R_{2}}\setminus\overline{B_{R_{1}}}$, $0<R_{1}<R_{2}$, and take $\Gamma_{0}=\partial B_{R_{2}}$, $\Gamma_{1}=\partial B_{R_{1}}$. The opposite choice of $\Gamma_{0}$ and $\Gamma_{1}$ would be possible, but this one seems to be more interesting than the other. Indeed it models a circular drumhead $B_{R_{2}}$ with a centered circular hole $\overline{B_{R_{1}}}$, in the example of the bassdrum.
@@ 2. Preliminaries
Given a Banach space $X$ we shall denote by $I$ the identity operator, by $X^{\prime}$ its dual and by $\langle\cdot,\cdot\rangle_{X}$ the duality product between them. When another Banach space $Y$ is given we shall denote by $\mathcal{L}(X,Y)$ the space of bounded linear operators between $X$ and $Y$.
For any $p\in(1,\infty)$ we shall denote by $p^{\prime}$ its Hölder conjugate, i.e. $11/p+1/p^{\prime}=$ 1, and for simplicity we shall denote by $\|\cdot\|_{p}$ the norms in $L^{p}(\Omega)$ and in $L^{p}(\Omega;\mathbb{R}^{N})$. Moreover, for any relatively open $\Gamma^{\prime}\subset\Gamma$, we shall denote $\|\cdot\|_{p,\Gamma^{\prime}}=\|\cdot\|_{L^{p}(\Gamma^{\prime})}$.
@@
In the sequel we shall systematically denote, without further mention, any relatively open subset of $\Gamma$ by $\Gamma^{\prime}$. Since $\Gamma$ is $C^{r}$ it inherits from $\mathbb{R}^{N}$ the structure of a Riemannian $C^{r}$ manifold, endowed with a $C^{r-1}$ Riemannian metric (see [41]), trivially restricting to $\Gamma^{\prime}$.
@@
The Riemannian gradient $\nabla_{\Gamma}$ is still defined, by density as $\nabla_{\Gamma}u=\sharp d_{\Gamma}\overline{u}$ for all $u\in H^{1}(\Gamma)$. It is well known, see for example [36, Chapter 3], that $H^{1}(\Gamma)$ can be equipped with the norm $\|\cdot\|_{H^{1}(\Gamma)}$, equivalent to the one in [21], given by
(23) $\|u\|_{H^{1}(\Gamma)}^{2}=\|u\|_{2,\Gamma}^{2}+\|\nabla_{\Gamma}u\|_{2,\Gamma}^{2},\quad\text{where}\quad\|\nabla_{\Gamma}u\|_{2,\Gamma}^{2}:=\int_{\Gamma}|\nabla_{\Gamma}u|_{\Gamma}^{2},$
@@
Trivially $H^{1}_{\Gamma_{0}}(\Gamma)$ is a closed subspace of $H^{1}(\Gamma)$. We shall endow it with the inherited norm. Due to the relevance of the space $H^{1}$ in our analysis it is useful to make some remarks on this space.
Since for all $u\in H^{1}_{\Gamma_{0}}(\Gamma)$ one has $\nabla_{\Gamma}u=0$ a.e. on $\Gamma\setminus\overline{\Gamma_{1}}$, and $\mathcal{H}^{N-1}(\overline{\Gamma_{0}}\cap\overline{\Gamma_{1}})=0$, we have
(25) $\|u\|_{H^{1}(\Gamma)}=\|u_{|\Gamma_{1}}\|_{H^{1}(\Gamma_{1})}\qquad\text{for all $u\in H^{1}_{\Gamma_{0}}(\Gamma)$,}$
suggesting the possibility of identifying $H^{1}_{\Gamma_{0}}(\Gamma)$ with $H^{1}(\Gamma_{1})$. Actually two different geometrical situations may occur.
@@
    When $\overline{\Gamma_{0}}\cap\overline{\Gamma_{1}}=\emptyset$, this case occurring when assumption (R) holds. Since the characteristic functions $\chi_{\Gamma_{i}}$, $i=0,1$ are $C^{r}$ on $\Gamma$, by identifying the elements of $H^{1}(\Gamma_{i})$, $i=0,1$, with their trivial extensions to $\Gamma$, we have the splitting
@@
    When $\overline{\Gamma_{0}}\cap\overline{\Gamma_{1}}\not=\emptyset$ the set $\Gamma_{0}$ is not relatively open on $\Gamma$ and (26)–(27) do not hold. Indeed, for example, $\chi_{\Gamma_{1}}\not\in H^{1}_{\Gamma_{0}}(\Gamma)$. In this case the elements of $H^{1}_{\Gamma_{0}}(\Gamma)$ “vanish” at the relative boundary $\partial\Gamma_{1}=\overline{\Gamma_{0}}\cap\overline{\Gamma_{1}}$ of $\Gamma_{1}$ on $\Gamma$, although such a notion can be made more precise only when $\partial\Gamma_{1}$ is regular enough. For example, if $\Gamma$ is smooth and $\overline{\Gamma_{1}}$ is a manifold with boundary $\partial\Gamma_{1}$, see [42, §5.1], then $H^{1}_{\Gamma_{0}}(\Gamma)$ is isometrically isomorphic to the space $H^{1}_{0}(\Gamma_{1}):=\overline{C^{\infty}_{c}(\Gamma_{1})}^{\|\cdot\|_{H^{1}(\Gamma_{1})}}$.
In the present paper we shall simultaneously deal with both cases A) and B) above up to the proof of Theorem 2, after which we shall restrict to the case A).
Since $g$, $g^{ij}$ are $C^{r-1}$ and $\Gamma$ is compact, when $\Gamma$ is $C^{2}$ formula (22) extends by density to $u\in W^{2,p}(\Gamma)$, $1<p<\infty$, so defining an operator $\Delta_{\Gamma}\in\mathcal{L}(W^{2,p}(\Gamma);L^{p}(\Gamma))$. This operator restricts, for $1<p<\infty$, and $s\in\mathbb{R}$, $11\leq s\leq r-$ 1, to $\Delta_{\Gamma}\in\mathcal{L}(W^{s+1,p}(\Gamma);W^{s-1,p}(\Gamma))$. Using again the compactness of $\Gamma$ and (22), integrating by parts and using a $C^{2}$ partition of the unity, one gets that
@@
Formula (28) motivates the definition of the operator $\Delta_{\Gamma}\in\mathcal{L}(W^{1,p}(\Gamma);W^{-1,p}(\Gamma))$ also when $\Gamma$ is merely $C^{1}$. Indeed, recalling that

### v2

@@ Abstract.
where $\Omega$ is a bounded open subset of $\mathbb{R}^{N}$ ($N\geq 2$) with a $C^{1}$ boundary $\Gamma=\Gamma_{0}\cup\Gamma_{1}$, $\Gamma_{0}\cap\Gamma_{1}=\emptyset$, $\Gamma_{1}$ being nonempty and relatively open on $\Gamma$. Moreover, $\mathcal{H}^{N-1}(\overline{\Gamma}_{0}\cap\overline{\Gamma}_{1})=0$ and $\mathcal{H}^{N-1}(\Gamma_{0})>0$. We prove that $L^{2}(\Omega)\times L^{2}(\Gamma_{1})$ admits a Hilbert basis constituted by eigenfunctions and we describe the behavior of the eigenvalues. Moreover, when $\Gamma$ is at least $C^{2}$ and $\overline{\Gamma}_{0}\cap\overline{\Gamma}_{1}=\emptyset$, we give several qualitative properties of the eigenfunctions.
@@ 1.1. Presentation of the problem and literature overwiew
where $\Omega$ is a bounded open subset of $\mathbb{R}^{N}$ ($N\geq 2$) with $C^{r}$ boundary $\Gamma$ (see [21]), with $r=1,2,\ldots,\infty$. Hence, when nothing is said, $r=1$. We also assume that $\Gamma=\Gamma_{0}\cup\Gamma_{1}$, $\Gamma_{0}\cap\Gamma_{1}=\emptyset$, $\Gamma_{1}$ being nonempty and relatively open on $\Gamma$ (or equivalently $\overline{\Gamma_{0}}=\Gamma_{0}$). Denoting by $\mathcal{H}^{N-1}$ the Hausdorff measure, we also assume that $\mathcal{H}^{N-1}(\overline{\Gamma}_{0}\cap\overline{\Gamma}_{1})=0$ and $\mathcal{H}^{N-1}(\Gamma_{0})>0$. These properties of $\Omega$, $\Gamma_{0}$ and $\Gamma_{1}$ will be assumed, without further comments, throughout the paper.
In problem (1) $\lambda$ is a real or complex parameter and we respectively denote by $\Delta$ and $\Delta_{\Gamma}$ the Laplace operator in $\Omega$ and the Laplace–Beltrami operator on $\Gamma$, while $\nu$ stands for the outward normal to $\Omega$. We shall look for eigenvalues and eigenfunctions of problem (1), that is for values of $\lambda$ for which (1) has a nontrivial (real or complex–valued) solution, i.e. an eigenfunction.
Problem (1) has been studied (as the particular case $K=0$, $\alpha=1$ and $\gamma=\omega$ in problem (1.2)) in [27], when $\Gamma_{0}=\emptyset$ and $\lambda,u$ are real. The study in [27] is motivated by several papers on the Allen–Cahn equation subject to a dynamic boundary condition, see [7, 9, 40]. Indeed finding a Hilbert basis of eigenfunctions allows to look for solutions of the evolution problem by using a Faedo–Galerkin scheme.
@@
As to the author’s knowledge problem (1) in the case $\Gamma_{0}\not=\emptyset$ has not yet been considered in the mathematical literature. The motivation for studying it originates from an evolution problem different than the one mentioned above. Indeed, it originates from the wave equation with hyperbolic boundary conditions. It is the evolutionary boundary value problem
@@
where $\Omega$, $\Gamma_{0}$ and $\Gamma_{1}$ are as above, $w=w(t,x)$, $t\in\mathbb{R}$, $x\in\overline{\Omega}$, $\Delta=\Delta_{x}$ and $\Delta_{\Gamma}$ denote the Laplacian and Laplace–Beltrami operators with respect to the space variable.
@@
where $u$ is nontrivial and real–valued. The function $w$ defined in (3) solves, at least formally, problem (2) if and only if $u$ solves problem (1) with $\lambda=\omega^{2}>0$.
Hence, in the analysis of problem (2), a deep understanding of the eigenvalue problem (1) would allow to find solutions by suing the Fourier method, provided one can find a complete system of eigenfunctions. Since in the analysis of problem (2) it is usefull to consider complex–valued solutions, in the sequel we shall consider complex–valued functions everywhere. The due attention will be given to find real eigenfunctions, so a reader only interested to the real case can simply ignore conjugation everywhere.
Since our study of problem (1) is motivated by problem (2), it is worth to give a brief overview of the literature dealing with it. Indeed, problems with hyperbolic or, more generally, kinetic boundary conditions, arise in several physical applications.
A one dimensional model was studied by several authors to describe transversal small oscillations of an elastic rod with a tip mass on one endpoint, while the other one is pinched. See [2, 10, 12, 22, 33].
A two dimensional model,  [^1] Although this model is two–dimensional, the title of the paper comes the model considered in [27], which is three dimensional, since the two models are formally strongly related. introduced in [20], more closely motivates problem (2). We shall briefly describe it. One considers a vibrating membrane of surface density $\mu>0$, subject to a tension $T>0$, both taken constant and normalized for simplicity. Moreover, $w=w(t,x)$, $t\in\mathbb{R}$, $x\in\Omega\subset\mathbb{R}^{2}$, denotes the vertical displacement from the rest state. After a standard linear approximation, $w$ satisfies the wave equation $w_{tt}-\Delta w=0$ in ×$\mathbb{R}\times\Omega$. Now one supposes that a part $\Gamma_{0}$ of the boundary is pinched, while the other part $\Gamma_{1}$ carries a linear mass density $m>0$ and it is subject to a linear tension $\tau>0$, both taken constant for simplicity.
A practical example of this situation is given by a drumhead with a hole in the interior having a thick border. This situation occurs in bass drums, as one can realize by looking at several pictures of them in the internet. During the paper we shall constantly refer to this motivating example as to the bass drum model.
@@
We also would like to point out that, when $\Gamma_{0}=\emptyset$ and $\Omega=\mathbb{R}^{N}_{+}$, problem (2) also shows up in Quantum Field Theory, see [53].
Several papers in the literature address the wave equation with kinetic boundary conditions. This fact is even more evident if one takes into account that, plugging the wave equation in (2) into the boundary condition, we can rewrite it as $-\Delta_{\Gamma}w+\partial_{\nu}w+\Delta w=0$, that is a generalized Wentzell (also spelled Ventcel) boundary condition. We refer to [15, 18, 34, 35, 51, 52], and also to the series of papers [46, 47, 48, 49] by the author. All the quoted papers deal with well–posedness issues for variously (linearly or nonlinearly) perturbed versions of (2).
@@
The analysis of the literature made above shows that the simple unperturbed problem (2), which well–posedness is rather standard, see for example [47], was object of a detailed study only in the case $\Gamma_{0}=\emptyset$ and $\Omega=\mathbb{R}^{N}_{+}$, see [53].
The aim of the present paper is to start such a study. In particular we shall analyze the eigenvalue problem (1). Our aim is to show that most of the well–known classical results concerning the homogeneous Dirichlet problem for the Helmholtz equation, that is the eigenvalue problem
@@
continue to hold for problem (1), in a suitably modified form. The parallelism will go further than one can expect. To illustrate the last assertion we shall present our main results in the sequel. Preliminarily we are now going to introduce some basic notation.
@@ 1.2. Functional spaces
where $\nabla_{\Gamma}$ denotes the Riemannian gradient on $\Gamma$, and $(\cdot,\cdot)_{|\Gamma}$ the unique Hermitian extension to the tangent bundle of the Riemannian metric on $\Gamma$, see § 2.2. We also denote $|\cdot|_{\Gamma}=(\cdot,\cdot)|_{\Gamma}^{1/2}$.
@@ 1.3. Main results
Moreover, an eigenvalue for (1) will be $\lambda\in\mathbb{C}$ for which (1) has a nontrivial weak solution, which is called an eigenfunction. Finally, for each eigenvalue $\lambda$, the subspace of weak solutions of (1) will be called the eigenspace associated to $\lambda$. The dimension of the eigenspace is called the eigenvalue’s multiplicity.
@@
Problem (1) has countably many eigenvalues, all of which are of finite multiplicity, constituting the set $\Lambda\subset\mathbb{R}$. By repeating each eigenvalue according to its multiplicity we can write $\Lambda=\{\lambda_{n},n\in\mathbb{N}\}$, where
(11) $0<\lambda_{1}\leq\lambda_{2}\leq\ldots\leq\lambda_{n}\leq\lambda_{n+1}\leq\cdots,\qquad\lambda_{n}\to\infty\quad\text{as $n\to\infty$.}$
Moreover there exists an orthogonal sequence $(u_{n})_{n}$ in $H^{1}$ such that each $u_{n}$ is a real–valued eigenfunction corresponding to $\lambda_{n}$, i.e. it is a nontrivial weak solution of
@@
Theorem 1 generalizes [27, Theorem 4.4], in the case $K=0$, $\alpha=1$ and $\gamma=\omega$, to $\Gamma_{0}\not=\emptyset$ and to the complex case.
@@
It extends to problem (1) well–known results concerning problem (4), see for example [17] or [14].
@@
Indeed we shall consider the case in which $\Gamma_{0}$ and $\Gamma_{1}$ are both relatively open on $\Gamma$, so $\Gamma$ has at least two connected components. We remark that, in the bass drum model described above, $\Gamma_{0}$ and $\Gamma_{1}$ are exactly these two connected components. Hence this assumption is rather natural. Moreover, we shall also assume $\Gamma$ to be at least $C^{2}$.
To avoid repeating these hypotheses several times we formalize them as the following assumption:
@@
both endowed with the norms inherited from the product spaces inside square brackets. Our third main result concerns the regularity of the eigenfunctions in Theorem 1. Beside its independent interest, this regularity result is an essential tool in pursuing the goal described above.
@@
Let assumption (R) hold and, for all $n\in\mathbb{N}$, let $u_{n}$ be a weak solution of (12). Then $u_{n}$ enjoys the following further regularity properties:
@@
The proof of Theorem 3 is based on appropriate regularity results for problem (13), see § 3 below, and on bootstrap arguments.
Our last main result brings the parallelism between problems (1) and (4) to a probably unexpected level. Indeed, despite of the difference between the two problems, their principal eigenspaces and eigenfunctions exhibit similar properties.
@@
Let assumption $(R)$ hold and let $\Omega$ be connected. Then the principal eigenvalue $\lambda_{1}$ in Theorem 1 is simple, i.e. $\lambda_{1}<\lambda_{2}$ in (11), and $u_{1}$ has constant sign in $\Omega\cup\Gamma_{1}$.
The most unexpected assertion in the last statement is that $u_{1}$ has constant sign also on $\Gamma_{1}$. The proof Theorem 4, indeed, uses classical arguments (see [17]) inside $\Omega$ and ad hoc arguments on $\Gamma_{1}$.
Finally, we would like to point out an interesting consequence of our main results in the radial case. Let $\Omega=B_{R_{2}}\setminus\overline{B_{R_{1}}}$, $0<R_{1}<R_{2}$, and take $\Gamma_{0}=\partial B_{R_{2}}$, $\Gamma_{1}=\partial B_{R_{1}}$. The opposite choice of $\Gamma_{0}$ and $\Gamma_{1}$ would be possible, but this one models a circular drumhead $B_{R_{2}}$ with a centered circular hole $\overline{B_{R_{1}}}$, in the bass drum model.
@@ 2. Preliminaries
Given a Banach space $X$, we shall denote by $I$ the identity operator, by $X^{\prime}$ its dual and by $\langle\cdot,\cdot\rangle_{X}$ the duality product between them. When another Banach space $Y$ is given we shall denote by $\mathcal{L}(X,Y)$ the space of bounded linear operators between $X$ and $Y$.
For any $p\in(1,\infty)$, we shall denote by $p^{\prime}$ its Hölder conjugate, i.e. $11/p+1/p^{\prime}=$ 1, and for simplicity we shall denote by $\|\cdot\|_{p}$ the norms in $L^{p}(\Omega)$ and in $L^{p}(\Omega;\mathbb{R}^{N})$. Moreover, for any relatively open $\Gamma^{\prime}\subset\Gamma$, we shall denote $\|\cdot\|_{p,\Gamma^{\prime}}=\|\cdot\|_{L^{p}(\Gamma^{\prime})}$.
@@
In the sequel we shall systematically denote by $\Gamma^{\prime}$, without further mention, any relatively open subset of $\Gamma$. Since $\Gamma$ is of class $C^{r}$, it inherits from $\mathbb{R}^{N}$ the structure of a Riemannian $C^{r}$ manifold, endowed with a $C^{r-1}$ Riemannian metric (see [41]), trivially restricting to $\Gamma^{\prime}$.
@@
The Riemannian gradient $\nabla_{\Gamma}$ is still defined, using a density argument, by $\nabla_{\Gamma}u=\sharp d_{\Gamma}\overline{u}$ for all $u\in H^{1}(\Gamma)$. It is well known, see for example [36, Chapter 3], that $H^{1}(\Gamma)$ can be equipped with the norm $\|\cdot\|_{H^{1}(\Gamma)}$, equivalent to the one introduced in [21], given by
(23) $\|u\|_{H^{1}(\Gamma)}^{2}=\|u\|_{2,\Gamma}^{2}+\|\nabla_{\Gamma}u\|_{2,\Gamma}^{2},\quad\text{where}\quad\|\nabla_{\Gamma}u\|_{2,\Gamma}^{2}:=\int_{\Gamma}|\nabla_{\Gamma}u|_{\Gamma}^{2}.$
@@
Trivially $H^{1}_{\Gamma_{0}}(\Gamma)$ is a closed subspace of $H^{1}(\Gamma)$. We shall endow it with the inherited norm. Due to the relevance of the space $H^{1}$ in our analysis it is useful to make some remarks on it.
Since, for all $u\in H^{1}_{\Gamma_{0}}(\Gamma)$, one has $\nabla_{\Gamma}u=0$ a.e. on $\Gamma\setminus\overline{\Gamma_{1}}$, and $\mathcal{H}^{N-1}(\overline{\Gamma_{0}}\cap\overline{\Gamma_{1}})=0$, we have
(25) $\|u\|_{H^{1}(\Gamma)}=\|u_{|\Gamma_{1}}\|_{H^{1}(\Gamma_{1})}\qquad\text{for all $u\in H^{1}_{\Gamma_{0}}(\Gamma)$.}$
Formula (25) suggests the possibility of identifying $H^{1}_{\Gamma_{0}}(\Gamma)$ with $H^{1}(\Gamma_{1})$. Actually two different geometrical situations may occur.
@@
    One can have $\overline{\Gamma_{0}}\cap\overline{\Gamma_{1}}=\emptyset$, this case occurring when assumption (R) holds. In it, since the characteristic functions $\chi_{\Gamma_{i}}$, $i=0,1$ are $C^{r}$ on $\Gamma$, by identifying the elements of $H^{1}(\Gamma_{i})$, $i=0,1$, with their trivial extensions to $\Gamma$, we have the splitting
@@
    One can have $\overline{\Gamma_{0}}\cap\overline{\Gamma_{1}}\not=\emptyset$. In this case the set $\Gamma_{0}$ is not relatively open on $\Gamma$ and (26)–(27) do not hold. Indeed, for example, $\chi_{\Gamma_{1}}\not\in H^{1}_{\Gamma_{0}}(\Gamma)$. In this case the elements of $H^{1}_{\Gamma_{0}}(\Gamma)$ “vanish” at the relative boundary $\partial\Gamma_{1}=\overline{\Gamma_{0}}\cap\overline{\Gamma_{1}}$ of $\Gamma_{1}$ on $\Gamma$, although such a notion can be made more precise only when $\partial\Gamma_{1}$ is regular enough. For example, if $\Gamma$ is smooth and $\overline{\Gamma_{1}}$ is a manifold with boundary $\partial\Gamma_{1}$, see [42, §5.1], then $H^{1}_{\Gamma_{0}}(\Gamma)$ is isometrically isomorphic to the space $H^{1}_{0}(\Gamma_{1}):=\overline{C^{\infty}_{c}(\Gamma_{1})}^{\|\cdot\|_{H^{1}(\Gamma_{1})}}$.
In the present paper we shall simultaneously deal with both cases A) and B) above up to the point where such a procedure is possible, that is up to the proof of Theorem 2. After it we shall restrict to the case A).
Since $g$, $g^{ij}$ are $C^{r-1}$ and $\Gamma$ is compact, when $\Gamma$ is $C^{2}$ formula (22) extends by density to $u\in W^{2,p}(\Gamma)$, $1<p<\infty$, so defining an operator $\Delta_{\Gamma}\in\mathcal{L}(W^{2,p}(\Gamma);L^{p}(\Gamma))$. This operator restricts, for $1<p<\infty$, and $s\in\mathbb{R}$, $11\leq s\leq r-$ 1, to $\Delta_{\Gamma}\in\mathcal{L}(W^{s+1,p}(\Gamma);W^{s-1,p}(\Gamma))$. Again using the compactness of $\Gamma$ and (22), integrating by parts and introducing a $C^{2}$ partition of the unity, one gets
@@
Formula (28) motivates the following definition of the operator
$\Delta_{\Gamma}\in\mathcal{L}(W^{1,p}(\Gamma);W^{-1,p}(\Gamma))$
also when $\Gamma$ is merely $C^{1}$. Indeed, recalling that